Solve the following equations using Gauss elimination and Gauss Jordon.


2x + 7y + z = 14

x + 3y - z = 2

x + 7y + 12z = 45


1
Expert's answer
2021-04-11T13:48:41-0400

Solution.

"2x + 7y + z = 14" ;

"x + 3y - z = 2" ;

"x + 7y + 12z = 45;"

"\begin{bmatrix}
 2 & 7&1&14 \\
 1 & 3&-1&2\\
1&7&12&45
\end{bmatrix}"

Find the pivot in the 1st column and swap the 2nd and the 1st rows

"\begin{bmatrix}
 1& 3&-1&2 \\
 2 & 7&1&14\\
1&7&12&45
\end{bmatrix}"

Eliminate the 1st column

"\begin{bmatrix}
 1& 3&-1&2 \\
 0 & 1&3&10\\
0&4&13&43
\end{bmatrix}"

Find the pivot in the 2nd column in the 2nd row

"\begin{bmatrix}
 1& 3&-1&2 \\
 0 & 1&3&10\\
0&4&13&43
\end{bmatrix}"

Eliminate the 2nd column

"\begin{bmatrix}
 1& 0&-10&-2 8\\
 0 & 1&3&10\\
0&0&1&3
\end{bmatrix}"

Find the pivot in the 3rd column in the 3rd row

"\begin{bmatrix}
 1& 0&-10&-2 8\\
 0 & 1&3&10\\
0&0&1&3
\end{bmatrix}"

Eliminate the 3rd column

"\begin{bmatrix}
 1& 0&0&2 \\
 0 & 1&0&1\\
0&0&1&3
\end{bmatrix}"

Solution set:

"x = 2;\n\ny = 1;\nz = 3."

Answer: "x = 2;\n\ny = 1;\nz = 3."